The difference between two numbers is -38. Two times the smaller number minus five times the larger number is -217. What are the numbers?
Answer by ichudov(499)

small number - big number = -38
2*(small number) - 5(big number) = -217
replace "small number" with S, "big number" with B.
S - B = -38
2S - 5B = -217.
This is a simple linear system, for which I have a solver that shows work, use it.